### v1

# for Large Language Models

Abstract
The capacity of Large Language Models (LLMs) to comprehend and reason over long contexts is pivotal for advancements in diverse fields. Yet, they still struggle with identifying relevant contexts and memory searching. To address this issue, we introduce Query-aware Inference for LLMs (Q-LLM), a system designed to process extensive sequences akin to human cognition. By focusing on memory data relevant to a given query, Q-LLM accurately captures pertinent information within a fixed window size and provides precise answers to queries. It requires no additional training and can be seamlessly integrated with any LLMs. Using LLaMA3 (QuickLLaMA), Q-LLM can read Harry Potter within 30 seconds and accurately answer related questions. On widely recognized benchmarks, Q-LLM improved performance by 7.17% compared to the current state-of-the-art on LLaMA3 and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by 7.0% on Mistral and achieved 100% accuracy on LLaMA3. Our code can be found in https://github.com/dvlab-research/Q-LLM.
[Uncaptioned image]QuickLLaMA: Query-aware Inference Acceleration
for Large Language Models
Jingyao Li1, Han Shi2, Xin Jiang2, Zhenguo Li2, Hong Xu1, Jiaya Jia1,3 1CUHK, 2Huawei Noah’s Ark Lab, 3SmartMore
1 Introduction
The ability to understand and reason over broad contexts has always been a long-term research focus of Large Language Models (LLMs) Dong et al. (2023). LLM-driven agents need to process ongoing information from external sources, which requires a strong ability to manage lengthy sequences Li et al. (2024b); Zheng et al. (2024); An ideal ChatBot assistant should be able to operate consistently over the content of conversations spanning recent days OpenAI et al. (2024). Other tasks such as summarizing and answering questions based on books, reports, and documents, as well as generating code at the repository level, also require the capability to handle long context sequences Bai et al. (2023); Zhang et al. (2024).
Figure 1: Examples of our QuickLLaMA-8B (1) reading long context containing 100K tokens, (2) reading our paper that has not be seen in the pretrained dataset, (3) retrieving value in long key-value pairs and (4) retrieving in Needle-in-a-Haystack task. More examples and comparisons with the SOTAs are provided in Appendix E.
Figure 2: This is an example from the $\infty$-Bench. Three questions were posed about the same long book: (1) Which among Annalisa, Seb, Peyton, and Gannonmarie is not Mrs. Bronwyn’s child? (2) What’s the name of the Bronwyns’ summer home? (3) Who among Mrs. Bronwyn, Mrs. Deandra, Rosemarie, and Cael is the final to perish? We present the score heatmap of the first 50 memory blocks. The methods used include (a) the consistent results from InfLLM for all three queries, and (b-d) the query-aware results from Q-LLM.
Yet, due to the challenges posed by unobserved extensive inputs Lin et al. (2024) and distracting, noisy contexts Liu et al. (2023); Tworkowski et al. (2024a), most LLMs that are pre-trained on sequences comprising a few thousand tokens struggle to generalize on longer sequences, resulting in unsatisfactory performance Press et al. (2022); Zhao et al. (2023). Some contemporary studies make use of sliding windows to disregard distant contexts, thereby ensuring that the length of the sequence do not surpass the LLMs’ maximum capacity Xiao et al. (2024b); Lin et al. (2024) and incorporate block-level context memory, which opts pertinent information from memory to disregard irrelevant disturbances Xiao et al. (2024a). However, the memory to be focused on should differ according to the specific query requirements. Yet, for distinct queries, InfLLM Xiao et al. (2024a) exhibits identical focal points when reading the long context, as shown in Fig. 2.
To address these challenges, we design Query-aware Inference for LLMs (Q-LLM), which processes extensive sequences in a manner similar to human cognition. Humans, when interpreting text, initially examine the question, and then seek the answer within the context, keeping the query in mind. This idea forms the foundation of our Query-aware Context Lookup strategy. Only memory data pertinent to the query is chosen for each computational step, disregarding unrelated distractions. As a result, LLMs can capture pertinent information within a fixed window size and provide precise answers to queries. Q-LLM doesn’t require extra training and can be seamlessly integrated with any LLMs.
We assess the performance of Q-LLM by utilizing LLaMA3-8B-inst AI@Meta (2024) (QuickLLaMA) and Mistral-7B-inst-v0.2 Jiang et al. (2023) as foundational models. These base models are pre-trained on sequences that do not exceed $8$K tokens. Instead, our QuickLLaMA can read Harry Potter with 100K tokens within half a minute on a single 800 GPU and accurately answer the questions, as shown in Fig. 1. We employ several widely recognized benchmarks, namely Longbench Bai et al. (2023), $\infty$-Bench Zhang et al. (2024), and Needle-in-a-Haystack Benchmark. Specifically, with a context window of 512, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral on the $\infty$-bench. In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by by 7.0% on Mistral and achieves 100% on LLaMA3. We have extended the input sequence to contexts of 1024K length, further demonstrating our model’s capability in handling extremely long sequences.
2 Related Works
The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al. (2023); Tay et al. (2022); Huang et al. (2023); Chen et al. (2023b). The approaches can generally be categorized into two main methods, our primary focus is on efficient context computation, which seeks to improve the computational efficiency of attention layers. This allows LLMs to make better use of limited resources when processing longer sequences. To ensure the completeness of the survey, in Appendix D, we provide an introduction to the other method, context length extrapolation. It aims to enable LLMs trained on short sequences to handle much longer sequences. We also mention the relevant work on memory-based models.
The computational and memory demands of LLM training often limit it to short sequences. Using LLMs directly on long sequences presents challenges such as out-of-domain issues and distractions from lengthy and noisy inputs Lin et al. (2024); Tworkowski et al. (2024a); Li et al. (2024d). As a result, context length extrapolation has emerged as a method to extend LLMs’ sequence length without additional training. Early approaches have designed new relative positional encoding mechanisms during pre-training Press et al. (2022); Tworkowski et al. (2024b). The following research has focused on the extensively adopted rotary position embedding (RoPE) Su et al. (2023), suggesting extending the length by interpolating positions to introduce non-integer positions Chen et al. (2024); Peng et al. (2023); Jin et al. (2024); Chen et al. (2023a). While these approaches can mitigate the out-of-domain issue stemming from unseen lengths, they fall short in addressing the distracting challenge posed by noisy context. In response to this limitation, Stream-LLM (Xiao et al., 2024b) and LM-Infinite(Lin et al., 2024) utilize the sliding window attention mechanism, discarding distant contexts to seamlessly process extremely long sequences. Additionally, InfLLM Xiao et al. (2024a) leverages a context memory to furnish LLMs with pertinent contextual information. These techniques, however, offer limited enhancement in capturing long-distance dependencies crucial for understanding lengthy texts. The model’s objective during long-text reading is inherently ambiguous, and it can become distracting when reading extensive articles. In this work, we introduce the Query-aware Context Lookup mechanism, enabling the model to effectively retrieve information relevant to the query from lengthy texts.
Figure 3: The illustration of our Q-LLM framework. The input from the memory context is partitioned into memory blocks, which are searched by Query-aware Context Lookup for query-related blocks. The current key-value cache comprises global tokens, query tokens, query-related blocks, and local tokens. Together, these form a new context window that, along with current tokens, is fed into the LLM.
3 Methods
In this section, we introduce the overall framework of Query-aware Inference for LLMs (Q-LLM) in Sec. 3.1, as depicted in Fig. 3. Then, we propose the preliminary memory block in Sec. 3.2 and our proposed Query-aware Context Lookup in Sec. 3.3.
3.1 Overall Framework
The primary challenges in expanding the context window of LLMs arise from issues related to out-of-domain and distractions, which are a result of the extensive and noisy context. To tackle these challenges, we follow prior studies, which implement the sliding window attention mechanism Xiao et al. (2024b); Lin et al. (2024) and the context memory module  Xiao et al. (2024a). Additionally, we propose the Query-aware Context Lookup strategy to find the query-related tokens from the context token. The past key-value vectors $\mathbf{P}=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{l_{P}}$ consist of four composers:
  1. 1.
    Global tokens $\mathbf{G}$, including system prompts and task description, etc.
  2. 2.
    Query tokens $\mathbf{Q}$, the query of the user.
  3. 3.
    Context tokens $\mathbf{C}$, the context stored in the context memory, consisting of multiple memory blocks.
  4. 4.
    Local tokens $\mathbf{L}$, the nearest tokens to the current token.
An example of these tokens in the prompt is shown in Fig. 4. We propose the Query-aware Context Lookup strategy to find the query-related tokens $\mathbf{R}$ from the context tokens $\mathbf{C}$:
$\mathbf{R}=\phi(\mathbf{H},\mathbf{C},\mathbf{Q}),$ (1)
where $\phi(\cdot)$ refers to the lookup operation of context memory. We will detail the strategy in Sec. 3.3. For each step, Q-LLM combines the global tokens, query tokens, query-related tokens, and local tokens to compose the current key-value cache.
$\mathbf{M}=\text{Concat}(\mathbf{G},\mathbf{Q},\mathbf{R},\mathbf{L}),$ (2)
Finally, the input parameters of the attention are:
$\mathbf{A_{q}}$ $=\mathbf{P_{q}}\mathbf{H},$ (3)
$\mathbf{A_{k}}$ $=\text{Concat}(\mathbf{M_{k}},\mathbf{P_{k}}\mathbf{H}),$
$\mathbf{A_{v}}$ $=\text{Concat}(\mathbf{M_{v}},\mathbf{P_{v}}\mathbf{H}),$
where $\mathbf{P_{q}}$, $\mathbf{P_{k}}$, and $\mathbf{P_{v}}$ are parameters in attention layers, $\mathbf{M_{k}}$ and $\mathbf{M_{v}}$ refer to the key and value vectors in the current key-value cache $\mathbf{M}$.
Figure 4: An example from Long-Bench. Global tokens include system prompts and task description. Query tokens represent the query of the user. Context tokens indicate the context stored in the context memory. We search query-related tokens from them, local tokens are the nearest tokens to the current token.
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
En.MC 26.64 27.95 26.63 29.69 31.00 30.13 33.19 33.19 30.13 30.57 33.62 34.50
Retrieve.PassKey 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0
Retrieve.Number 3.39 3.39 99.73 100.0 3.39 3.39 99.83 99.83 3.39 3.39 44.58 40.00
Code.Debug 31.22 32.74 31.02 31.22 35.79 37.56 38.58 38.58 35.79 32.74 35.03 34.52
Math.Find 17.71 17.14 24.77 25.93 16.57 17.43 27.71 28.86 15.43 16.29 28.29 28.29
Retrieve.KV 0.20 0.20 13.18 32.40 0.40 0.40 32.60 47.80 1.00 1.00 62.94 73.00
Average 13.76 14.14 49.22 53.21 15.09 15.38 55.32 58.04 14.86 14.56 50.74 51.72
(a) $\infty$-Bench (214K tokens)
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
NarrativeQA 8.80 9.77 11.80 12.04 9.44 10.19 15.61 15.95 12.44 13.37 18.75 20.14
Qasper 9.19 9.45 16.13 15.45 10.93 10.73 19.15 19.24 14.58 15.04 20.78 19.97
MultiFieldQA 25.38 26.03 38.43 41.35 27.82 27.76 42.65 43.71 32.29 32.02 43.74 44.83
HotpotQA 19.68 20.46 28.19 27.32 22.16 21.91 32.47 34.47 23.21 23.70 34.66 36.53
2WikiMQA 12.27 12.63 13.70 15.22 13.85 13.32 16.14 16.57 17.13 17.51 17.99 19.97
Musique 6.45 6.55 11.38 12.99 7.91 7.64 14.74 15.27 9.81 11.30 12.16 17.05
GovReport 22.50 22.40 29.64 28.46 24.79 24.90 30.18 29.82 27.07 27.12 30.26 29.75
QMSum 18.74 18.93 21.55 21.69 19.23 19.19 22.03 22.27 19.67 19.52 21.55 22.36
MultiNews 23.23 23.28 25.19 24.95 25.51 25.41 26.15 26.39 25.95 26.10 26.71 26.84
TREC 38.00 39.50 45.50 47.50 30.50 29.00 48.00 49.50 31.00 28.25 47.50 48.25
TriviaQA 79.68 80.54 82.02 82.20 85.06 84.27 83.20 84.56 88.06 87.08 82.81 84.49
SAMSum 35.30 34.58 36.65 37.18 36.05 35.09 38.20 38.12 36.30 36.21 37.91 38.25
PassageRetrieval 4.40 5.54 13.29 25.04 7.92 7.92 25.67 31.04 18.21 18.46 40.29 49.67
LCC 50.06 51.59 50.14 48.61 50.94 53.27 50.83 51.10 52.20 54.95 54.59 54.52
RepoBench-P 47.38 48.04 42.92 41.32 48.85 51.31 41.75 43.21 47.36 47.60 45.08 45.90
Average 25.18 25.59 29.30 30.22 26.40 26.44 31.80 32.70 28.56 28.77 33.55 35.06
(b) Long-Bench (31K tokens)
Table 1: The results comparison based on Mistral-7B-inst-v0.2 Jiang et al. (2023). The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K. Our results are highlighted in teal and best results are indicated in bold.
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
En.MC 37.12 34.93 37.77 40.17 37.12 34.93 37.99 40.17 36.24 37.12 33.19 34.50
Retrieve.PassKey 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0
Retrieve.Number 3.39 3.39 96.61 98.98 3.39 3.39 40.68 41.19 3.39 3.39 28.64 27.63
Code.Debug 22.59 22.59 22.59 22.59 22.59 22.59 23.10 23.86 24.11 22.84 22.59 23.10
Math.Find 20.86 19.71 29.23 30.70 20.86 19.71 32.29 31.14 18.00 16.86 26.86 27.37
Retrieve.KV 0.80 0.80 24.40 61.20 0.80 0.80 57.20 70.80 1.80 1.80 80.80 84.00
Average 14.69 14.14 51.77 58.94 14.69 14.14 48.54 51.19 14.49 14.23 48.68 49.43
(a) $\infty$-Bench (214K tokens)
Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
Context Window 512 1K 2K
NarrativeQA 14.50 14.56 19.28 19.29 14.50 14.56 19.90 20.50 16.47 15.12 19.41 25.60
Qasper 21.06 20.72 26.08 26.58 21.06 20.72 32.35 31.47 32.01 31.72 41.27 39.12
MultiFieldQA 25.66 25.79 36.01 40.12 25.66 25.79 41.46 46.44 31.63 30.99 45.89 48.30
HotpotQA 31.95 32.84 41.42 42.34 31.95 32.84 43.75 49.15 34.73 35.26 44.97 49.91
2WikiMQA 24.72 24.28 28.44 29.63 24.72 24.28 30.83 31.53 29.22 30.59 36.27 39.63
Musique 12.72 13.62 17.48 18.75 12.72 13.62 21.26 23.95 13.50 13.64 19.73 25.03
GovReport 26.25 25.93 29.26 26.83 26.25 25.93 30.44 28.73 27.84 27.83 30.68 29.80
QMSum 19.38 19.42 19.10 19.02 19.38 19.42 20.30 20.62 19.91 20.14 21.36 22.23
MultiNews 26.42 26.57 26.63 25.22 26.42 26.57 27.46 26.85 27.36 27.37 27.87 27.85
TriviaQA 82.46 82.47 80.81 77.65 82.46 82.47 85.11 85.04 88.07 87.35 88.03 87.70
SAMSum 38.28 37.91 38.73 38.83 38.28 37.91 39.59 40.40 36.93 35.97 34.86 34.97
PassageRetrieval 13.75 13.25 19.50 35.00 13.75 13.25 58.75 69.00 23.50 23.50 85.25 88.00
LCC 54.17 53.88 53.24 52.39 56.32 53.88 60.23 60.43 60.42 58.15 58.17 58.37
RepoBench-P 60.71 60.59 58.33 57.04 62.81 60.59 60.86 60.51 64.95 62.97 62.01 61.04
Average 32.29 32.27 35.31 36.34 32.59 32.27 40.88 42.47 36.18 35.76 43.98 45.54
(b) Long-Bench (31K tokens)
Table 2: The results comparison based on LLaMA3-8B-inst AI@Meta (2024). The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K. Our results are highlighted in teal and best results are indicated in bold.
(a) Mistral-7B-inst-v0.2
(b) LLaMA3-8B-inst
Figure 5: The comparison of performance in the Needle-in-a-Haystack task. The horizontal axis represents the document’s length (the haystack), whereas the vertical axis specifies the location of a brief sentence (the needle) within the document, ranging from 1K to 128K tokens. For instance, a 50% value implies that the needle is situated halfway through the document. A red cell indicates the language model’s inability to recall the needle’s information, while a green cell denotes successful recall by the model.
3.2 Memory Block
In light of the local semantic coherence in extended sequences, referring to previous studies Xiao et al. (2024a), we perform a memory lookup at the block level. We segment the context tokens $\mathbf{C}$ into multiple memory blocks, each containing $l_{b}$ tokens. We then select $n_{r}$ tokens that have the highest representative scores to represent the block. For the $i$-th token, the representative score is calculated as
$r_{i}=\frac{1}{l_{L}}\sum_{j=1}^{l_{L}}\mathbf{q}_{i+j}\cdot\mathbf{k}_{i},$ (4)
where $l_{L}$ is the length of local token, $\mathbf{q}_{i+j}$ is the query vector for the $(i+j)$-th token and $\mathbf{k}_{i}$ is the key vector for the $i$-th token. The score $r_{i}$ intuitively measures the importance of the $i$-th token within its local window, demonstrating its influence on other tokens in the same window.
3.3 Query-aware Context Lookup
When humans read and comprehend text, they first read the question and then search for the answer within the context with the question in mind. For instance, in Fig. 4, when reading a novel with the question "What nationality is Ruth Anvory?", we can quickly locate the query-related memory context, which is "…that she was American". Building on this concept, we introduce Query-aware Context Lookup, a simple but efficient lookup strategy. Our defined criterion for search is to locate tokens relevant to the query. We propose the relevance score between a memory block $\mathbf{B}$ and query tokens $\mathbf{Q}$ as follows:
$s(\mathbf{B},\mathbf{Q})=\sum_{i=1}^{l_{Q}}\sum_{j=1}^{r_{k}}\mathbf{Q_{q}}_{i}\cdot\mathbf{B_{k}}_{j}^{r},$ (5)
where $l_{Q}$ is the length of query tokens. $\mathbf{Q}_{q_{i}}$ is the $i$-th query vector of $\mathbf{Q}$ and $\mathbf{B_{k}}_{j}^{r}$ is the $j$-th representative key vector of $\mathbf{B}$. The score $s(\mathbf{B},\mathbf{Q})$ is independent of the current token $\mathbf{H}$, therefore, it only needs to be calculated once during inference.
On the other hand, the importance of different memory blocks is influenced by varying current tokens Xiao et al. (2024a). Therefore, the relevance score with the current token is also a criterion for selecting a memory block. The relevance score between a memory block $\mathbf{B}$ and current tokens $\mathbf{H}$ is defined as:
$s(\mathbf{B},\mathbf{H})=\sum_{i=1}^{l_{H}}\sum_{j=1}^{r_{k}}\mathbf{H_{q}}_{i}\cdot\mathbf{B_{k}}_{j}^{r},$ (6)
where $l_{H}$ is the length of current tokens. $\mathbf{H}_{q_{i}}$ is the $i$-th query vector of $\mathbf{H}$ and $\mathbf{B_{k}}_{j}^{r}$ is the $j$-th representative key vector of $\mathbf{B}$. The final memory block score is thus composed of these two components:
$s(\mathbf{B})=s(\mathbf{B},\mathbf{H})+\beta s(\mathbf{B},\mathbf{Q}),$ (7)
where $\beta$ represents the balancing factor. We opt to store the $n_{b}$ memory blocks with the highest scores in the current key-value cache. In terms of intuition, $s(\mathbf{B},\mathbf{Q})$ and $s(\mathbf{B},\mathbf{H})$ respectively express the search for the memory blocks related to the query tokens $\mathbf{Q}$ and the current tokens $\mathbf{H}$. Ablation experiments in Sec. 4.3 validate that $\beta\geq 1$, indicating that the selection of queries is more crucial to the memory block. This aligns with our initial motivation. More methodology details are in Appendix A.
4 Experiments
In this section, we conduct experiments utilizing Mistral-7B-inst-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (infllm) Xiao et al. (2024a). We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). More configuration details are in Appendix C. We use the same prompts for all methods to ensure a fair comparison, as detailed in Appendix F.
Figure 6: Ablation of query weight $\beta$.
4.1 Long-Bench and $\infty$-Bench
In this section, we utilize representative tasks from two widely-recognized long document benchmarks, $\infty$-Bench Zhang et al. (2024) and Long-Bench Bai et al. (2023). These benchmarks encompass a variety of tasks including question answering, summarization, few-shot learning, context retrieval, mathematical computing, code completion, etc. The $95$% quantile for sequence lengths in $\infty$-Bench and Long-Bench reaches $214$K and $31$K tokens. The outcomes based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are detailed in Tab. 1 and Tab. 2 respectively. The following observations can be made from the results: (1) Our approach shows considerable enhancement in performance when compared to models utilizing the sliding window mechanism (StreamingLLM and LM-Infinite) across benchmarks and context window lengths. This suggests that the context memory in Q-LLM can effectively provide LLMs with appropriate contextual data, facilitating efficient comprehension and reasoning on long sequences. (2) Our method also exhibits a significant performance uplift when compared to models with other lookup mechanisms (InfLLM). This implies that previous methods still struggle to extract valid information from noisy contexts. Our proposed Query-aware Context Lookup, however, can purposefully use the query to find relevant information in the long context. (3) Our technique is particularly beneficial in scenarios with longer input contexts and relatively smaller available context windows, as observed in comparisons across different benchmarks and context budgets. For instance, with a context window of 512 on the $\infty$-bench, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral. This illustrates our model’s superiority in infinite stream scenarios. (4) Our model’s relative improvement is more noticeable on LLaMA3 when compared to other models. This is because superior models can more effectively utilize query information to precisely locate relevant information in the long context. This paves the way for our model to further enhance its inference capabilities with the development of superior LLMs in future.
4.2 Needle in a Haystack
Needle-in-a-Haystack Kamradt (2023) is a widely used benchmark to evaluate if models can effectively utilize extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (needle) that is randomly positioned within a document that could be as long as 128K (haystack). The results for methods based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are shown in Fig. 5a and Fig. 5b. The context window size is 512. Our method accurately locates the needle within the haystack across 1K to 128K tokens. Specifically, Q-LLM improved upon InfLLM by 7.0% on Mistral and achieved 100% on LLaMA3. Notably, even though InfLLM performs well in tasks like Retrieve.PassKey in $\infty$-Bench, which is less likely to generate illusions, it does not perform as well in the Needle-in-a-Haystack Bench. InfLLM generates illusions from the long book, confusing the needle with irrelevant content in the long context, leading to subpar performance, as depicted in Fig. 1. In contrast, our query-aware method effectively locates and alleviates this confusion by focusing on query-related information, resulting in superior performance.
4.3 Ablation Experiments
To further substantiate the efficacy of Query-aware Context Lookup, we carry out ablation studies in this section, with results displayed in Fig. 6. The performance of Mistral-7B-inst-v0.2 and LLaMA3-8B-inst exhibits a trend of initial increase followed by a decrease as the query weight $\beta$ escalates. We select the $\beta$ at the peak point as the experimental setup, choosing $\beta=1$ for Mistral and $\beta=4$ for LLaMA3. The maximum $\beta$ corresponds to LLaMA3, indicating that superior models can more effectively harness the query information to retrieve valid data from the context memory accurately. This aligns with our previous observations and lays the groundwork for our model to further bolster its inference prowess with the evolution of large language models. Further exploration of the primary elements within the context memory is in Appendix B.
4.4 Scaling up
In this sub-section, we’re evaluating Q-LLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K. We use the Retrieve.KV task from the $\infty$-Bench to test its ability to discern context in extensive sequences. This task requires LLMs to identify a value from a key and a dictionary, essentially locating pertinent information within long sequences. For each length, 50 instances are created for assessment. The base model used is Mistral-7B-inst-v0.2. The outcomes are displayed in Fig. 7. According to the results, even when the context length is scaled to $1024$ thousand tokens, Q-LLM consistently performs at a level significantly above the current state-of-the-art. This confirms Q-LLM’s ability to accurately recognize long-distance dependencies for effective long-sequence reasoning.
Figure 7: The results for Q-LLM and InfLLM on sequences with extremely lengthy sequence lengths.
5 Conclusion
In this study, we focused on the significant challenges faced by LLMs in processing and reasoning over extensive contexts. We introduced Q-LLM, an approach inspired by human cognitive processes, which focuses on relevant memory data and effectively bypasses context input clutter. Our method does not require additional training and can be seamlessly integrated with any LLM. Through comprehensive evaluations using the LLaMA and Mistral models on the Longbench, $\infty$-Bench, and Needle-in-a-Haystack benchmarks, Q-LLM demonstrated a marked improvement over the current SOTA. Moreover, our QuickLLaMA can read 100K tokens within 30 seconds. The empirical results validate Q-LLM’s ability to capture long-range dependencies and manage vast contexts efficiently, paving the way for enhanced performance in LLM-driven tasks that require long-sequence reasoning.
Limitations
While Q-LLM demonstrates promising improvements over the current SOTA methods in various benchmarks, there are still some limitations. For instance, the system’s performance relies on the limited window size, which could lead to potential information loss when dealing with highly complex contexts. Future research should address these limitations and explore the potential of Q-LLM in a broader range of tasks and contexts.
Broader Impact
The advancements made by Q-LLM in understanding and reasoning over broad contexts, a long-term research focus of Large Language Models (LLMs), could have profound implications across various fields. Given its ability to manage lengthy sequences,Q-LLM’s potential to operate consistently over the content of conversations spanning recent days could make ChatBot assistants more effective and user-friendly. Tasks such as summarizing and answering questions based on books, reports, and documents, as well as generating code at the repository level, could also be improved with the ability to handle long context sequences. However, it is crucial to recognize that the benefits of Q-LLM also come with potential risks. The system’s ability to process and understand extensive contexts could be misused for nefarious purposes, such as creating deepfakes or other forms of misinformation.
Appendix A Methodology Details
In this section, we introduce our methodology details.
A.1 Chunks
Given the constraints of GPU memory, we do not encode the input sequence at once Xiao et al. (2024a); instead, we process it in chunks and generate output on a token-by-token basis. In each computational step, the inputs are composed of past key-value vectors $\mathbf{P}=\{(\mathbf{k}_{i},\mathbf{v}_{i})\}_{i=1}^{l_{P}}$ and current token hidden vectors $\mathbf{H}=\{\mathbf{h}_{i}\}_{i=1}^{l_{H}}$. When encoding, $l_{H}$ is equivalent to the chunk size, while during decoding, $l_{H}$ is equal to one.
A.2 Positional Encoding
Traditional LLM training typically utilizes a limited set of positional encodings, which can face difficulties with out-of-domain distribution when extended to process longer sequences Xiao et al. (2024a). Furthermore, in Q-LLM, the current key-value cache consists of several discontinuous text blocks. Assigning continuous positional encodings to these blocks could create mismatches and confuse the model. Consequently, drawing upon previous studies Raffel et al. (2023); Su (2023); Xiao et al. (2024a), we assign identical positional encodings to all tokens exceeding the local window size. More precisely, we set the distance between tokens in context memory units and the current tokens as $l_{L}$.
A.3 Cache Management
In order to process exceedingly lengthy sequence streams and encapsulate the semantic relevance present in extended contexts with LLMs Xiao et al. (2024a), it’s necessary to maintain all memory units and reference them at every computational stage. Given that most units are seldom used, we adopt an offloading strategy, which stores most memory units in CPU memory. Only the tokens and memory units essential for current operations are kept in GPU memory. Furthermore, due to the semantic continuity in long sequences where neighboring tokens often necessitate similar memory units, we assign a cache area in GPU memory, governed by a least recently used policy. This method enables efficient encoding of exceptionally long sequences using finite GPU memory. Moreover, for extremely long sequences, the representative tokens for each unit can be offloaded to CPU memory, forming an effective k-nearest-neighbor index, thereby further diminishing computational complexity.
Appendix B Further Exploration
(a) Representative Tokens
(b) Block Size
(c) Number of Blocks
Figure 8: Further exploration, investigating the influence of the context memory with varying numbers of representative tokens, selected memory blocks, and memory block sizes, respectively.
Q-LLM leverages context memory to retrieve pertinent data. We delve deeper into the influence of primary elements within the context memory. Results are presented in  Fig. 8. Conduct experiments on Mistral-7B-inst-v0.2 using the default parameters with a context window length of 1024.
Number of Representative Tokens. Q-LLM divides key-value vectors into memory blocks and picks a few representative tokens from each block to act as the segment’s representation. The capacity of these tokens to symbolize the entire segment semantically directly impacts the model’s efficacy. We run tests with the different quantity of representative tokens. The outcomes are displayed in Fig. 8a. We note an upward trend in model performance as the number of tokens increases, suggesting that a larger token count can better capture the semantic essence of memory segments. However, when the token count hits $8$, a slight performance dip is seen in HotpotQA. This drop can be traced back to the inclusion of semantically unrelated tokens in the segment representations. Future work could enhance model performance by developing more efficient and potent segment representations.
Memory Block Size. Each memory block should ideally represent a consistent semantic block. Oversized blocks can obstruct precise lookup, while undersized ones can escalate the computational cost of memory lookup. We test Q-LLM with different block sizes while maintaining a total context length of $1024$. The outcomes are displayed in  Fig. 8b. It’s evident that the best block size changes according to the task due to the differing characteristics of input sequences. For instance, in NarrativeQA, a semantically complete paragraph forms a semantic block, while in Retrieve.Number, a single number does. Using heuristic rules to segment context can result in less-than-optimal performance. Thus, investigating how to dynamically segment context is a vital future research direction.
Number of Memory blocks. The chosen blocks are employed to supply relevant context to LLMs. We run tests with different number of blocks. From Fig. 8c, we see that as the number of chosen blocks rises, there’s a significant improvement in model performance. This is because more blocks mean a higher recall rate of relevant content. However, a larger block count also increases the time needed for memory scheduling and the computational time for attention. Hence, advancing lookup accuracy continues to be a key area for enhancing the efficiency of LLMs.
Appendix C Configuration Details
Context Window Local Tokens Block Size Block Num
512 256 64 4
1024 512 64 8
2048 1024 128 8
Table 3: The parameters for different context window length, including number of local tokens, memory block size and number of memory blocks.
Datasets. We utilize representative tasks from following widely-recognized long document benchmarks:
(1) $\infty$-Bench Zhang et al. (2024) and LongBench Bai et al. (2023). Given that our base models are primarily pre-trained on English corpora, we employ English datasets for the evaluation. These benchmarks encompass a variety of tasks such as question answering, summarization, few-shot learning, context retrieval, mathematical computing, and code completion. The average document length in $\infty$-Bench is $145.1$K tokens, and in LongBench, it is $12.8$K tokens. The $95$% quantile for sequence lengths in these benchmarks reaches $214$K and $31$K tokens respectively, which significantly exceeds the maximum length of the base models.
(2) Needle-in-a-Haystack Kamradt (2023), a widely used benchmark to evaluate if models can effectively use extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (referred to as the needle) that is randomly positioned within a document that could be as long as 128K (referred to as the haystack).
Baseline Methods. Our goal is to enable LLMs trained with limited sequence lengths to comprehend extremely long sequences without additional training. For this purpose, we use Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. Mistral-7B-Instruct-v0.2 is initially pre-trained with a maximum sequence length of $8$K tokens and subsequently fine-tuned with a maximum sequence length of $32$K tokens. LLaMA3-8B-inst is fine-tuned from LLaMA3-8B AI@Meta (2024) with an $8$K token maximum length. We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (infllm) Xiao et al. (2024a). The same prompts are utilized for all methods and detailed in Appendix F.
Implementation Details. Following the settings of previous works Xiao et al. (2024a), the number of representative tokens, $r_{k}$, is set to 4. The number of initial tokens is set to 128 for LM-Infinite, StreamingLLM, InfLLM, and Q-LLM, to cover system prompts and task descriptions. We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). The local tokens, block size, and number of blocks for InfLLM and Q-LLM are detailed in Tab. 3.
Appendix D Detailed Related Works
The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al. (2023); Tay et al. (2022); Huang et al. (2023); Chen et al. (2023b); Yu et al. (2023). The approaches can generally be categorized into two main methods, our primary focus is on the efficient context computation, which seeks to improve the computational efficiency of attention layers. This allows LLMs to make better use of limited resources when processing longer sequences. To ensure the completeness of the survey, in this section, we provide an introduction to the other method, context length extrapolation, which aims to enable LLMs trained on short sequences to handle much longer sequences, as well as the relevant work on memory-based models.
D.1 Efficient Context Computation
The computational complexity of attention layers, which grows quadratically, is a significant bottleneck restricting LLMs’ capability to handle lengthy sequences. Consequently, numerous researchers have devised efficient attention mechanisms, including sparse attention Zaheer et al. (2021); Beltagy et al. (2020); Child et al. (2019); Ainslie et al. (2020); Zhao et al. (2019), approximate attention computations using kernel functions Kitaev et al. (2020); Wang et al. (2020); Katharopoulos et al. (2020), and replacing attention layers with state-space models of linear complexity Gu et al. (2022); Gu and Dao (2023). These approaches necessitate architectural modifications, requiring retraining of the models. Concurrently, many scholars have tackled this challenge from an infrastructural angle by optimizing the memory usage of attention computations to mitigate the computational resource requirements of the model Dao et al. (2022); Dao (2023); Hong et al. (2024); Shazeer (2019); Kwon et al. (2023). Given the training-free nature of our method, it can be seamlessly integrated to further expedite LLM inference.
D.2 Memory-based Approaches
Memory networks have been extensively researched for decades and have demonstrated effectiveness in enhancing models with additional information storage capabilities Graves et al. (2014); Weston et al. (2015); Sukhbaatar et al. (2015); Miller et al. (2016). With the rise of pre-trained models, memory layers have gradually found application in the training stage of recurrent transformer layers, enabling models to recursively process long sequences Dai et al. (2019); Rae et al. (2020); Khandelwal et al. (2020); Wu et al. (2022); Bertsch et al. (2023). These approaches segment sequences, encoding each segments individually, and utilize memory to retain context information from preceding segments. Although these methodologies share conceptual similarities with Q-LLM, they necessitate architectural modifications and are typically incorporated during the pre-training phase. In contrast, our objective is to explore the intrinsic properties of LLMs and introduce a training-free Query-aware Context Lookup mechanism for long-text comprehension.
Figure 9: Examples of our QuickLLaMA-8B summarizing multiple papers Li et al. (2023c, 2024a, a, 2024b).
Appendix E More Examples
In this section, we provide examples of our QuickLLaMA-8B.
  1. 1.
    Figure 9 summarizes multiple papers.
  2. 2.
    Figure 10 summarizes paper series and discovering their connection and comparison.
  3. 3.
    Figure 11 reviews our paper according to ACL review guideline.
  4. 4.
    Figure 12 reads long novel Journey to the West comparing with current SOTAs.
Figure 10: Examples of our QuickLLaMA-8B summarizing paper series and discovering their connection and comparison Li et al. (2023b, 2024c).
Figure 11: Examples of our QuickLLaMA-8B reviewing our paper according to ACL review guideline.
Figure 12: Examples of our QuickLLaMA-8B reading long novel Journey to the West comparing with current SOTAs
Appendix F Prompts
We utilize the same prompts for all experiments in the main paper.
  1. 1.
    The prompts for Long-Bench are in Tabs. 4 and 5.
  2. 2.
    The prompts for $\infty$-Bench are in Tab. 6.
  3. 3.
    The prompts for Needle-in-a-Haystack Bench are in Tab. 7.
Dataset Prompt
NarrativeQA You are given a story, which can be either a novel or a movie script, and a question. Answer the question asconcisely as you can, using a single phrase if possible. Do not provide any explanation.
Question: {input}
Story: {context}
Now, answer the question based on the story asconcisely as you can, using a single phrase if possible. Do not provide any explanation.
Question: {input}
Answer:
Qasper You are given a scientific article and a question. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write "unanswerable". If the question is a yes/no question, answer "yes", "no", or "unanswerable". Do not provide any explanation.
Question: {input}
Article: {context}
Answer the question based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write "unanswerable". If the question is a yes/no question, answer "yes", "no", or "unanswerable". Do not provide any explanation.
Question: {input}
Answer:
MultiFieldQA Read the following text and answer briefly.
Question: {input}
{context}
Now, answer the following question based on the above text, only give me the answer and do not output any other words.
Question: {input}
Answer:
HotpotQA Answer the question based on the given passages. Only give me the answer and do not output any other words.
The following are given passages.
Question: {input}
{context}
Answer the question based on the given passages. Only give me the answer and do not output any other words.
Question: {input}
Answer:
2WikiMQA Answer the question based on the given passages. Only give me the answer and do not output any other words.
The following are given passages.
Question: {input}
{context}
Answer the question based on the given passages. Only give me the answer and do not output any other words.
Question: {input}
Answer:
Musique Answer the question based on the given passages. Only give me the answer and do not output any other words.
The following are given passages.
Question: {input}
{context}
Answer the question based on the given passages. Only give me the answer and do not output any other words.
Question: {input}
Answer:
GovReport You are given a report by a government agency. Write a one-page summary of the report.
Report:
{context}
Now, write a one-page summary of the report.
Summary:
Table 4: The prompt for each dataset in Long-Bench Bai et al. (2023). Yellow highlights indicate the query, while a gray background represents the long context.
Dataset Prompt
QMSum You are given a meeting transcript and a query containing a question or instruction. Answer the query in one or more sentences.
Query: {input}
Transcript:
{context}
Now, answer the query based on the above meeting transcript in one or more sentences.
Query: {input}
Answer:
MultiNews You are given several news passages. Write a one-page summary of all news.
News:
{context}
Now, write a one-page summary of all the news.
Summary:
TREC Please determine the type of the question below.
{input}
Here are some examples of questions.
{context}
Now please determine the type of the question below.
{input}
TriviaQA Answer the question based on the given passage. Only give me the answer and do not output any other words.
{input}
The following are some examples.
{context}
Now answer the question based on the given passage. Only give me the answer and do not output any other words.
{input}
SAMSum Summarize the dialogue into a few short sentences.
{input}
The following are some examples.
{context}
{input}
PassageRetrieval Here are 30 paragraphs from Wikipedia, along with an abstract. Please determine which paragraph the abstract is from.
Abstract:
{input}
Paragraphs:
{context}
Please enter the number of the paragraph that the abstract is from. The answer format must be like "Paragraph 1", "Paragraph 2", etc.
The answer is:
LCC Please complete the code given below .
{context}Next line of code:
RepoBench-P Please complete the code given below .
{context} {input}
Next line of code:
Table 5: The prompt for each dataset in Long-Bench Bai et al. (2023) (continued). Yellow highlights indicate the query, while a gray background represents the long context.
Dataset Prompt
En.MC Read the book and answer the question.Only one of the following options is correct, tell me the answer using one single letter (A, B, C, or D). Don’t say anything else.
Question: {question}
A. {OPTIONA}
B. {OPTIONB}
C. {OPTIONC}
D. {OPTIOND}
{context}
Question: {question}
Only one of the following options is correct, tell me the answer using one single letter (A, B, C, or D). Don’t say anything else.
A. {OPTIONA}
B. {OPTIONB}
C. {OPTIONC}
D. {OPTIOND}
Retrieve.PassKey There is an important info hidden inside a lot of irrelevant text. Find and memorize it: {input}
{context}
{input}
Retrieve.Number There is an important info hidden inside a lot of irrelevant text. Find and memorize it: {input}
{context}
{input}
Code.Debug There is ONLY ONE function in the large project that is deliberately made to include an obvious error. Please find the function that contains the most obvious errors. I will give you four options to narrow your scope. You can inspect the options and think. Eventually, tell me the answer using one single letter (A, B, C, or D).
Which funtion has deliberate error?
A. {OPTIONA}
B. {OPTIONB}
C. {OPTIONC}
D. {OPTIOND}
{context}
Which funtion has deliberate error?
A. {OPTIONA}
B. {OPTIONB}
C. {OPTIONC}
D. {OPTIOND}
Give me your answer for the function that has the deliberate and obvious error in A, B, C, or D. Your answer MUST be chosen from one of the four options without any explanation. If you cannot determine answers accurately, you also MUST provide the answer you think is most likely. Absolutely do not say you do not know or you need more information.
Math.Find {prefix}
{context}
{input}
Retrieve.KV Extract the value corresponding to the specified key {key} in the JSON object below.
{context}
{input}
Table 6: The prompt for each dataset in $\infty$-Bench Zhang et al. (2024) (continued). Yellow highlights indicate the query, while a gray background represents the long context.
Dataset Prompt
Needle in a Haystack Based on the content of the book, Question: What is the best thing to do in San Francisco? <book> <context> </book>.
Table 7: The prompt for each dataset in Needle-in-a-Haystack Benchmark Kamradt (2023). Yellow highlights indicate the query, while a gray background represents the long context.

### v2

@@
 Yet, they still struggle with identifying relevant contexts and memory searching.
-To address this issue, we introduce Query-aware Inference for LLMs (Q-LLM), a system designed to process extensive sequences akin to human cognition.
-By focusing on memory data relevant to a given query, Q-LLM accurately captures pertinent information within a fixed window size and provides precise answers to queries.
+To address this issue, we introduce Query-aware Inference for LLMs (QLLM), a system designed to process extensive sequences akin to human cognition.
+By focusing on memory data relevant to a given query, QLLM accurately captures pertinent information within a fixed window size and provides precise answers to queries.
 It requires no additional training and can be seamlessly integrated with any LLMs.
-Using LLaMA3 (QuickLLaMA), Q-LLM can read Harry Potter within 30 seconds and accurately answer related questions.
-On widely recognized benchmarks, Q-LLM improved performance by 7.17% compared to the current state-of-the-art on LLaMA3 and by 3.26% on Mistral on the $\infty$-bench.
-In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by 7.0% on Mistral and achieved 100% accuracy on LLaMA3.
-Our code can be found in https://github.com/dvlab-research/Q-LLM.
+Using LLaMA3 (QuickLLaMA), QLLM can read Harry Potter within 30 seconds and accurately answer related questions.
+On widely recognized benchmarks, QLLM improved performance by 7.17% compared to the current state-of-the-art on LLaMA3 and by 3.26% on Mistral on the $\infty$-bench.
+In the Needle-in-a-Haystack and BABILong task, QLLM improved upon the current SOTA by 7.0% and 6.1%.
+Our code can be found in https://github.com/dvlab-research/Q-LLM
 [Uncaptioned image]QuickLLaMA: Query-aware Inference Acceleration
@@
 Figure 1: Examples of our QuickLLaMA-8B (1) reading long context containing 100K tokens, (2) reading our paper that has not be seen in the pretrained dataset, (3) retrieving value in long key-value pairs and (4) retrieving in Needle-in-a-Haystack task.
-More examples and comparisons with the SOTAs are provided in Appendix E.
+More examples and comparisons with the SOTAs are provided in Appendix D.
 Figure 2: This is an example from the $\infty$-Bench.
@@
 We present the score heatmap of the first 50 memory blocks.
-The methods used include (a) the consistent results from InfLLM for all three queries, and (b-d) the query-aware results from Q-LLM.
+The methods used include (a) the consistent results from InfLLM for all three queries, and (b-d) the query-aware results from QLLM.
 Yet, due to the challenges posed by unobserved extensive inputs Lin et al.
@@
 2.
-To address these challenges, we design Query-aware Inference for LLMs (Q-LLM), which processes extensive sequences in a manner similar to human cognition.
+To address these challenges, we design Query-aware Inference for LLMs (QLLM), which processes extensive sequences in a manner similar to human cognition.
 Humans, when interpreting text, initially examine the question, and then seek the answer within the context, keeping the query in mind.
@@
 As a result, LLMs can capture pertinent information within a fixed window size and provide precise answers to queries.
-Q-LLM doesn’t require extra training and can be seamlessly integrated with any LLMs.
-We assess the performance of Q-LLM by utilizing LLaMA3-8B-inst AI@Meta (2024) (QuickLLaMA) and Mistral-7B-inst-v0.2 Jiang et al.
+QLLM doesn’t require extra training and can be seamlessly integrated with any LLMs.
+We assess the performance of QLLM by utilizing LLaMA3-8B-inst AI@Meta (2024) (QuickLLaMA) and Mistral-7B-inst-v0.2 Jiang et al.
 (2023) as foundational models.
 These base models are pre-trained on sequences that do not exceed $8$K tokens.
-Instead, our QuickLLaMA can read Harry Potter with 100K tokens within half a minute on a single 800 GPU and accurately answer the questions, as shown in Fig.
-1.
+Instead, our QuickLLaMA can read Harry Potter with 100K tokens within half a minute on a single A800 GPU and accurately answer the questions, as shown in Fig.
+7.
 We employ several widely recognized benchmarks, namely Longbench Bai et al.
@@
 (2024), and Needle-in-a-Haystack Benchmark.
-Specifically, with a context window of 512, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral on the $\infty$-bench.
-In the Needle-in-a-Haystack task, Q-LLM improved upon the current SOTA by by 7.0% on Mistral and achieves 100% on LLaMA3.
-We have extended the input sequence to contexts of 1024K length, further demonstrating our model’s capability in handling extremely long sequences.
+Specifically, with a context window of 512, QLLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral on the $\infty$-bench.
+In the Needle-in-a-Haystack task, QLLM improved upon the current SOTA by by 7.0% on Mistral and achieves 100% on LLaMA3.
+In the BABILong task, QLLM improved upon the current SOTA by 6.1%.
+We have extended the input sequence to contexts of 1048K length, further demonstrating our model’s capability in handling extremely long sequences.
+Figure 3: The illustration of our QLLM framework.
+The input from the memory context is partitioned into memory blocks, which are searched by Query-aware Context Lookup for query-related blocks.
+The current key-value cache comprises global tokens, query tokens, query-related blocks, and local tokens.
+Together, these form a new context window that, along with current tokens, is fed into the LLM.
 2 Related Works
-The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al.
-(2023); Tay et al.
-(2022); Huang et al.
-(2023); Chen et al.
-(2023b).
-The approaches can generally be categorized into two main methods, our primary focus is on efficient context computation, which seeks to improve the computational efficiency of attention layers.
-This allows LLMs to make better use of limited resources when processing longer sequences.
-To ensure the completeness of the survey, in Appendix D, we provide an introduction to the other method, context length extrapolation.
-It aims to enable LLMs trained on short sequences to handle much longer sequences.
-We also mention the relevant work on memory-based models.
+Efficient Context Computation.
 The computational and memory demands of LLM training often limit it to short sequences.
@@
 (2024); Chen et al.
-(2023a).
-While these approaches can mitigate the out-of-domain issue stemming from unseen lengths, they fall short in addressing the distracting challenge posed by noisy context.
-In response to this limitation, Stream-LLM (Xiao et al., 2024b) and LM-Infinite(Lin et al., 2024) utilize the sliding window attention mechanism, discarding distant contexts to seamlessly process extremely long sequences.
+(2023).
+To process extremely long sequences, Stream-LLM (Xiao et al., 2024b) and LM-Infinite(Lin et al., 2024) utilize the sliding window attention mechanism and discard distant contexts.
 Additionally, InfLLM Xiao et al.
 (2024a) leverages a context memory to furnish LLMs with pertinent contextual information.
-These techniques, however, offer limited enhancement in capturing long-distance dependencies crucial for understanding lengthy texts.
-The model’s objective during long-text reading is inherently ambiguous, and it can become distracting when reading extensive articles.
+Yet, the objective of these models during long-text reading is inherently ambiguous, and it can become distracting when reading extensive articles.
 In this work, we introduce the Query-aware Context Lookup mechanism, enabling the model to effectively retrieve information relevant to the query from lengthy texts.
-Figure 3: The illustration of our Q-LLM framework.
-The input from the memory context is partitioned into memory blocks, which are searched by Query-aware Context Lookup for query-related blocks.
-The current key-value cache comprises global tokens, query tokens, query-related blocks, and local tokens.
-Together, these form a new context window that, along with current tokens, is fed into the LLM.
+Context Length Extrapolation.
+The computational complexity of attention layers, which grows quadratically, is a significant bottleneck restricting LLMs’ capability to handle lengthy sequences.
+Consequently, numerous researchers have devised efficient attention mechanisms, including sparse attention Zaheer et al.
+(2021); Beltagy et al.
+(2020); Child et al.
+(2019); Ainslie et al.
+(2020); Zhao et al.
+(2019), approximate attention computations using kernel functions Kitaev et al.
+(2020); Wang et al.
+(2020); Katharopoulos et al.
+(2020), and replacing attention layers with state-space models of linear complexity Gu et al.
+(2022); Gu and Dao (2023).
+These approaches necessitate architectural modifications, requiring retraining of the models.
+Concurrently, many scholars have tackled this challenge from an infrastructural angle by optimizing the memory usage of attention computations to mitigate the computational resource requirements of the model Dao et al.
+(2022); Dao (2023); Hong et al.
+(2024); Shazeer (2019); Kwon et al.
+(2023).
+Given the training-free nature of our method, it can be seamlessly integrated to further expedite LLM inference.
+Memory-based Approaches.
+Memory networks have been extensively researched for decades and have demonstrated effectiveness in enhancing models with additional information storage capabilities Graves et al.
+(2014); Weston et al.
+(2015); Sukhbaatar et al.
+(2015); Miller et al.
+(2016).
+With the rise of pre-trained models, memory layers have gradually found application in the training stage of recurrent transformer layers, enabling models to recursively process long sequences Dai et al.
+(2019); Rae et al.
+(2020); Khandelwal et al.
+(2020); Wu et al.
+(2022); Bertsch et al.
+(2023).
+These approaches segment sequences, encoding each segments individually, and utilize memory to retain context information from preceding segments.
+Yet, they necessitate architectural modifications and are typically incorporated during the pre-training phase.
+In contrast, our objective is to explore the intrinsic properties of LLMs and introduce a training-free Query-aware Context Lookup mechanism for long-text comprehension.
 3 Methods
-In this section, we introduce the overall framework of Query-aware Inference for LLMs (Q-LLM) in Sec.
+In this section, we introduce the overall framework of Query-aware Inference for LLMs (QLLM) in Sec.
 3.1, as depicted in Fig.
@@
     Local tokens $\mathbf{L}$, the nearest tokens to the current token.
-An example of these tokens in the prompt is shown in Fig. 4. We propose the Query-aware Context Lookup strategy to find the query-related tokens $\mathbf{R}$ from the context tokens $\mathbf{C}$:
+An example of these tokens in the prompt is shown in Fig.
+4.
+Given that all memory blocks are necessary to be maintained and most of them and seldom used, we adopt an offloading strategy, which stores most memory blocks in CPU memory.
+More details are in Sec.
+A.3.
+We propose the Query-aware Context Lookup strategy to find the query-related tokens $\mathbf{R}$ from the context tokens $\mathbf{C}$:
 $\mathbf{R}=\phi(\mathbf{H},\mathbf{C},\mathbf{Q}),$ (1)
@@
 3.3.
-For each step, Q-LLM combines the global tokens, query tokens, query-related tokens, and local tokens to compose the current key-value cache.
+For each step, QLLM combines the global tokens, query tokens, query-related tokens, and local tokens to compose the current key-value cache.
 $\mathbf{M}=\text{Concat}(\mathbf{G},\mathbf{Q},\mathbf{R},\mathbf{L}),$ (2)
@@
 We search query-related tokens from them, local tokens are the nearest tokens to the current token.
-Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
+Method Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
 Context Window 512 1K 2K
@@
 (a) $\infty$-Bench (214K tokens)
-Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
+Method Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
 Context Window 512 1K 2K
@@
 (b) Long-Bench (31K tokens)
-Table 1: The results comparison based on Mistral-7B-inst-v0.2 Jiang et al.
-(2023).
-The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K.
-Our results are highlighted in teal and best results are indicated in bold.
-Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
-Context Window 512 1K 2K
-En.MC 37.12 34.93 37.77 40.17 37.12 34.93 37.99 40.17 36.24 37.12 33.19 34.50
-Retrieve.PassKey 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0
-Retrieve.Number 3.39 3.39 96.61 98.98 3.39 3.39 40.68 41.19 3.39 3.39 28.64 27.63
-Code.Debug 22.59 22.59 22.59 22.59 22.59 22.59 23.10 23.86 24.11 22.84 22.59 23.10
-Math.Find 20.86 19.71 29.23 30.70 20.86 19.71 32.29 31.14 18.00 16.86 26.86 27.37
-Retrieve.KV 0.80 0.80 24.40 61.20 0.80 0.80 57.20 70.80 1.80 1.80 80.80 84.00
-Average 14.69 14.14 51.77 58.94 14.69 14.14 48.54 51.19 14.49 14.23 48.68 49.43
+Table 1: The results comparison based on Mistral-7B-inst-v0.2 Jiang et al. (2023). Our results are highlighted in teal and best results are indicated in bold.
+Method LLaMA Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
+Context Window -1048K 512 1K 2K
+En.MC 31.0 37.12 34.93 37.77 40.17 37.12 34.93 37.99 40.17 36.24 37.12 33.19 34.50
+Retrieve.PassKey 6.78 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0 3.40 3.40 100.0 100.0
+Retrieve.Number 6.78 3.39 3.39 96.61 98.98 3.39 3.39 40.68 41.19 3.39 3.39 28.64 27.63
+Code.Debug 22.59 22.59 22.59 22.59 22.59 22.59 22.59 23.10 23.86 24.11 22.84 22.59 23.10
+Math.Find 34.29 20.86 19.71 29.23 30.70 20.86 19.71 32.29 31.14 18.00 16.86 26.86 27.37
+Retrieve.KV 6.2 0.80 0.80 24.40 61.20 0.80 0.80 57.20 70.80 1.80 1.80 80.80 84.00
+Average 17.94 14.69 14.14 51.77 58.94 14.69 14.14 48.54 51.19 14.49 14.23 48.68 49.43
 (a) $\infty$-Bench (214K tokens)
-Method Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM Infinite Stream Infllm Q-LLM
-Context Window 512 1K 2K
-NarrativeQA 14.50 14.56 19.28 19.29 14.50 14.56 19.90 20.50 16.47 15.12 19.41 25.60
-Qasper 21.06 20.72 26.08 26.58 21.06 20.72 32.35 31.47 32.01 31.72 41.27 39.12
-MultiFieldQA 25.66 25.79 36.01 40.12 25.66 25.79 41.46 46.44 31.63 30.99 45.89 48.30
-HotpotQA 31.95 32.84 41.42 42.34 31.95 32.84 43.75 49.15 34.73 35.26 44.97 49.91
-2WikiMQA 24.72 24.28 28.44 29.63 24.72 24.28 30.83 31.53 29.22 30.59 36.27 39.63
-Musique 12.72 13.62 17.48 18.75 12.72 13.62 21.26 23.95 13.50 13.64 19.73 25.03
-GovReport 26.25 25.93 29.26 26.83 26.25 25.93 30.44 28.73 27.84 27.83 30.68 29.80
-QMSum 19.38 19.42 19.10 19.02 19.38 19.42 20.30 20.62 19.91 20.14 21.36 22.23
-MultiNews 26.42 26.57 26.63 25.22 26.42 26.57 27.46 26.85 27.36 27.37 27.87 27.85
-TriviaQA 82.46 82.47 80.81 77.65 82.46 82.47 85.11 85.04 88.07 87.35 88.03 87.70
-SAMSum 38.28 37.91 38.73 38.83 38.28 37.91 39.59 40.40 36.93 35.97 34.86 34.97
-PassageRetrieval 13.75 13.25 19.50 35.00 13.75 13.25 58.75 69.00 23.50 23.50 85.25 88.00
-LCC 54.17 53.88 53.24 52.39 56.32 53.88 60.23 60.43 60.42 58.15 58.17 58.37
-RepoBench-P 60.71 60.59 58.33 57.04 62.81 60.59 60.86 60.51 64.95 62.97 62.01 61.04
-Average 32.29 32.27 35.31 36.34 32.59 32.27 40.88 42.47 36.18 35.76 43.98 45.54
+Method LLaMA Infinite Stream ILM QLLM Infinite Stream ILM QLLM Infinite Stream ILM QLLM
+Context Window -1048K 512 1K 2K
+NarrativeQA 23.78 14.50 14.56 19.28 19.29 14.50 14.56 19.90 20.50 16.47 15.12 19.41 25.60
+Qasper 21.22 21.06 20.72 26.08 26.58 21.06 20.72 32.35 31.47 32.01 31.72 41.27 39.12
+MultiFieldQA 39.89 25.66 25.79 36.01 40.12 25.66 25.79 41.46 46.44 31.63 30.99 45.89 48.30
+HotpotQA 17.16 31.95 32.84 41.42 42.34 31.95 32.84 43.75 49.15 34.73 35.26 44.97 49.91
+2WikiMQA 18.11 24.72 24.28 28.44 29.63 24.72 24.28 30.83 31.53 29.22 30.59 36.27 39.63
+Musique 10.39 12.72 13.62 17.48 18.75 12.72 13.62 21.26 23.95 13.50 13.64 19.73 25.03
+GovReport 33.76 26.25 25.93 29.26 26.83 26.25 25.93 30.44 28.73 27.84 27.83 30.68 29.80
+QMSum 23.38 19.38 19.42 19.10 19.02 19.38 19.42 20.30 20.62 19.91 20.14 21.36 22.23
+MultiNews 27.68 26.42 26.57 26.63 25.22 26.42 26.57 27.46 26.85 27.36 27.37 27.87 27.85
+TriviaQA 87.76 82.46 82.47 80.81 77.65 82.46 82.47 85.11 85.04 88.07 87.35 88.03 87.70
+SAMSum 41.89 38.28 37.91 38.73 38.83 38.28 37.91 39.59 40.40 36.93 35.97 34.86 34.97
+PassageRetrieval 51.83 13.75 13.25 19.50 35.00 13.75 13.25 58.75 69.00 23.50 23.50 85.25 88.00
+LCC 43.79 54.17 53.88 53.24 52.39 56.32 53.88 60.23 60.43 60.42 58.15 58.17 58.37
+RepoBench-P 46.11 60.71 60.59 58.33 57.04 62.81 60.59 60.86 60.51 64.95 62.97 62.01 61.04
+Average 34.77 32.29 32.27 35.31 36.34 32.59 32.27 40.88 42.47 36.18 35.76 43.98 45.54
 (b) Long-Bench (31K tokens)
-Table 2: The results comparison based on LLaMA3-8B-inst AI@Meta (2024).
-The $95$% quantile for text lengths in these two benchmarks is $214$K and $31$K.
-Our results are highlighted in teal and best results are indicated in bold.
-(a) Mistral-7B-inst-v0.2
-(b) LLaMA3-8B-inst
-Figure 5: The comparison of performance in the Needle-in-a-Haystack task.
-The horizontal axis represents the document’s length (the haystack), whereas the vertical axis specifies the location of a brief sentence (the needle) within the document, ranging from 1K to 128K tokens.
-For instance, a 50% value implies that the needle is situated halfway through the document.
-A red cell indicates the language model’s inability to recall the needle’s information, while a green cell denotes successful recall by the model.
+Table 2: The results comparison based on LLaMA3-8B-inst AI@Meta (2024). Our results are highlighted in teal and best results are indicated in bold.
 3.2 Memory Block
@@
 More methodology details are in Appendix A.
+Figure 5: The comparison of performance in the Needle-in-a-Haystack task.
+The horizontal axis represents the document’s length (the haystack), whereas the vertical axis specifies the location of a brief sentence (the needle) within the document, ranging from 1K to 128K tokens.
+A red cell indicates the language model’s inability to recall the needle’s information, while a green cell denotes successful recall by the model.
 4 Experiments
@@
 (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models.
-We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al.
+We compare our methods with LLaMA3-8B-inst-1048K (LLaMA-1048K) Gradient (2024) and other competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al.
 (2024), StreamingLLM (Stream) Xiao et al.
@@
 More configuration details are in Appendix C.
-We use the same prompts for all methods to ensure a fair comparison, as detailed in Appendix F.
-Figure 6: Ablation of query weight $\beta$.
+Note that we add the queries before the context to ensure the baselines also have query-aware capabilities, as detailed in Appendix E.
+Figure 6: (a) Memory and (b) Time consumption of different methods as tokens increases.
+(c) Ablation of query weight $\beta$.
+(d) The results of methods on sequences with extremely lengthy sequence lengths.
 4.1 Long-Bench and $\infty$-Bench
@@
 (2023).
-These benchmarks encompass a variety of tasks including question answering, summarization, few-shot learning, context retrieval, mathematical computing, code completion, etc.
 The $95$% quantile for sequence lengths in $\infty$-Bench and Long-Bench reaches $214$K and $31$K tokens.
@@
 2 respectively.
-The following observations can be made from the results: (1) Our approach shows considerable enhancement in performance when compared to models utilizing the sliding window mechanism (StreamingLLM and LM-Infinite) across benchmarks and context window lengths.
-This suggests that the context memory in Q-LLM can effectively provide LLMs with appropriate contextual data, facilitating efficient comprehension and reasoning on long sequences.
+The following observations can be made from the results: (1) Our approach shows considerable enhancement in performance when compared to base models (LLaMA3-8B-inst-1048K ) and that utilizing the sliding window mechanism (StreamingLLM and LM-Infinite) across benchmarks and context window lengths.
+This suggests that the context memory in QLLM can effectively provide LLMs with appropriate contextual data, facilitating efficient comprehension and reasoning on long sequences.
 (2) Our method also exhibits a significant performance uplift when compared to models with other lookup mechanisms (InfLLM).
@@
 (3) Our technique is particularly beneficial in scenarios with longer input contexts and relatively smaller available context windows, as observed in comparisons across different benchmarks and context budgets.
-For instance, with a context window of 512 on the $\infty$-bench, Q-LLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral.
+For instance, with a context window of 512 on the $\infty$-bench, QLLM improved by 7.17% compared to the current SOTA on LLaMA3, and by 3.26% on Mistral.
 This illustrates our model’s superiority in infinite stream scenarios.
@@
 This is because superior models can more effectively utilize query information to precisely locate relevant information in the long context.
-This paves the way for our model to further enhance its inference capabilities with the development of superior LLMs in future.
 4.2 Needle in a Haystack
@@
 The results for methods based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are shown in Fig.
-5a and Fig.
-5b.
+5.
 The context window size is 512.
 Our method accurately locates the needle within the haystack across 1K to 128K tokens.
-Specifically, Q-LLM improved upon InfLLM by 7.0% on Mistral and achieved 100% on LLaMA3.
-Notably, even though InfLLM performs well in tasks like Retrieve.PassKey in $\infty$-Bench, which is less likely to generate illusions, it does not perform as well in the Needle-in-a-Haystack Bench.
-InfLLM generates illusions from the long book, confusing the needle with irrelevant content in the long context, leading to subpar performance, as depicted in Fig.
-1.
-In contrast, our query-aware method effectively locates and alleviates this confusion by focusing on query-related information, resulting in superior performance.
+Specifically, QLLM improved upon InfLLM by 7.0% on Mistral and achieved 100% on LLaMA3.
+BABILong Kuratov et al.
+(2024) is a challenging benchmark for evaluating the performance of models in processing arbitrarily long documents with distributed facts.
+We conducted experiments on the BABILong based on LLaMA3-8B-inst.
+At a window length of 1024, our method improved by 6.1% compared to InfLLM.
+Figure 7: Average accuracy over QA1-QA5 tasks from BABILong. The horizontal and vertical axes represent the context and window length.
+4.3 Time and Memory Consumption
+In Fig.
+7(a) and (b), we compare the time and memory consumption of different input tokens across methods.
+QLLM and InfLLM are comparable in terms of efficiency and memory usage.
+The time consumed by InfLLM and QLLM increases almost linearly with the number of input tokens, requiring only 25.6 seconds and 22.3GB of memory to process 100k tokens.
+In contrast, LLaMA3-8B-inst-1048K shows a rapid increase in time and memory consumption with the number of input tokens and cannot handle 16k tokens on a single A800 GPU (maximum memory of 80GB).
+The context length for InfLLM and QLLM is 2048.
 4.4 Ablation Experiments
 To further substantiate the efficacy of Query-aware Context Lookup, we carry out ablation studies in this section, with results displayed in Fig.
-6.
+7(c).
 The performance of Mistral-7B-inst-v0.2 and LLaMA3-8B-inst exhibits a trend of initial increase followed by a decrease as the query weight $\beta$ escalates.
 We select the $\beta$ at the peak point as the experimental setup, choosing $\beta=1$ for Mistral and $\beta=4$ for LLaMA3.
-The maximum $\beta$ corresponds to LLaMA3, indicating that superior models can more effectively harness the query information to retrieve valid data from the context memory accurately.
-This aligns with our previous observations and lays the groundwork for our model to further bolster its inference prowess with the evolution of large language models.
 Further exploration of the primary elements within the context memory is in Appendix B.
 4.5 Scaling up
-In this sub-section, we’re evaluating Q-LLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K.
-We use the Retrieve.KV task from the $\infty$-Bench to test its ability to discern context in extensive sequences.
-This task requires LLMs to identify a value from a key and a dictionary, essentially locating pertinent information within long sequences.
-For each length, 50 instances are created for assessment.
-The base model used is Mistral-7B-inst-v0.2.
+In this sub-section, we’re evaluating QLLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K.
+The base model used is Mistral-7B-inst-v0.2 and the task is Retrieve.KV from $\infty$-Bench.
 The outcomes are displayed in Fig.
-7.
-According to the results, even when the context length is scaled to $1024$ thousand tokens, Q-LLM consistently performs at a level significantly above the current state-of-the-art.
-This confirms Q-LLM’s ability to accurately recognize long-distance dependencies for effective long-sequence reasoning.
-Figure 7: The results for Q-LLM and InfLLM on sequences with extremely lengthy sequence lengths.
+7(d).
+According to the results, even when the context length is scaled to $1024$ thousand tokens, QLLM consistently performs at a level significantly above the current state-of-the-art.
+This confirms QLLM’s ability to accurately recognize long-distance dependencies for effective long-sequence reasoning.
 5 Conclusion
 In this study, we focused on the significant challenges faced by LLMs in processing and reasoning over extensive contexts.
-We introduced Q-LLM, an approach inspired by human cognitive processes, which focuses on relevant memory data and effectively bypasses context input clutter.
+We introduced QLLM, an approach inspired by human cognitive processes, which focuses on relevant memory data and effectively bypasses context input clutter.
 Our method does not require additional training and can be seamlessly integrated with any LLM.
-Through comprehensive evaluations using the LLaMA and Mistral models on the Longbench, $\infty$-Bench, and Needle-in-a-Haystack benchmarks, Q-LLM demonstrated a marked improvement over the current SOTA.
+Through comprehensive evaluations using the LLaMA and Mistral models on the Longbench, $\infty$-Bench, and Needle-in-a-Haystack benchmarks, QLLM demonstrated a marked improvement over the current SOTA.
 Moreover, our QuickLLaMA can read 100K tokens within 30 seconds.
-The empirical results validate Q-LLM’s ability to capture long-range dependencies and manage vast contexts efficiently, paving the way for enhanced performance in LLM-driven tasks that require long-sequence reasoning.
+The empirical results validate QLLM’s ability to capture long-range dependencies and manage vast contexts efficiently, paving the way for enhanced performance in LLM-driven tasks that require long-sequence reasoning.
 Limitations
-While Q-LLM demonstrates promising improvements over the current SOTA methods in various benchmarks, there are still some limitations.
+While QLLM demonstrates promising improvements over the current SOTA methods in various benchmarks, there are still some limitations.
 For instance, the system’s performance relies on the limited window size, which could lead to potential information loss when dealing with highly complex contexts.
-Future research should address these limitations and explore the potential of Q-LLM in a broader range of tasks and contexts.
+Future research should address these limitations and explore the potential of QLLM in a broader range of tasks and contexts.
 Broader Impact
-The advancements made by Q-LLM in understanding and reasoning over broad contexts, a long-term research focus of Large Language Models (LLMs), could have profound implications across various fields.
-Given its ability to manage lengthy sequences,Q-LLM’s potential to operate consistently over the content of conversations spanning recent days could make ChatBot assistants more effective and user-friendly.
+The advancements made by QLLM in understanding and reasoning over broad contexts, a long-term research focus of Large Language Models (LLMs), could have profound implications across various fields.
+Given its ability to manage lengthy sequences,QLLM’s potential to operate consistently over the content of conversations spanning recent days could make ChatBot assistants more effective and user-friendly.
 Tasks such as summarizing and answering questions based on books, reports, and documents, as well as generating code at the repository level, could also be improved with the ability to handle long context sequences.
-However, it is crucial to recognize that the benefits of Q-LLM also come with potential risks.
+However, it is crucial to recognize that the benefits of QLLM also come with potential risks.
 The system’s ability to process and understand extensive contexts could be misused for nefarious purposes, such as creating deepfakes or other forms of misinformation.
@@
 (2024a).
-Furthermore, in Q-LLM, the current key-value cache consists of several discontinuous text blocks.
+Furthermore, in QLLM, the current key-value cache consists of several discontinuous text blocks.
 Assigning continuous positional encodings to these blocks could create mismatches and confuse the model.
@@
 (2024a), we assign identical positional encodings to all tokens exceeding the local window size.
-More precisely, we set the distance between tokens in context memory units and the current tokens as $l_{L}$.
+More precisely, we set the distance between tokens in context memory blocks and the current tokens as $l_{L}$.
 A.3 Cache Management
-In order to process exceedingly lengthy sequence streams and encapsulate the semantic relevance present in extended contexts with LLMs Xiao et al.
-(2024a), it’s necessary to maintain all memory units and reference them at every computational stage.
-Given that most units are seldom used, we adopt an offloading strategy, which stores most memory units in CPU memory.
-Only the tokens and memory units essential for current operations are kept in GPU memory.
-Furthermore, due to the semantic continuity in long sequences where neighboring tokens often necessitate similar memory units, we assign a cache area in GPU memory, governed by a least recently used policy.
+In order to process exceedingly lengthy sequence streams and encapsulate the semantic relevance with LLMs Xiao et al.
+(2024a), it’s necessary to maintain all memory blocks and reference them at every computational stage.
+Given that most blocks are seldom used, we adopt an offloading strategy, which stores most memory blocks in CPU memory.
+Only the tokens and memory blocks essential for current operations are kept in GPU memory.
+Furthermore, due to the semantic continuity in long sequences where neighboring tokens often necessitate similar memory blocks, we assign a cache area in GPU memory, governed by a least recently used policy.
 This method enables efficient encoding of exceptionally long sequences using finite GPU memory.
-Moreover, for extremely long sequences, the representative tokens for each unit can be offloaded to CPU memory, forming an effective k-nearest-neighbor index, thereby further diminishing computational complexity.
+Moreover, for extremely long sequences, the representative tokens for each block can be offloaded to CPU memory, forming an effective k-nearest-neighbor index, thereby further diminishing computational complexity.
 Appendix B Further Exploration
@@
 Figure 8: Further exploration, investigating the influence of the context memory with varying numbers of representative tokens, selected memory blocks, and memory block sizes, respectively.
-Q-LLM leverages context memory to retrieve pertinent data.
+QLLM leverages context memory to retrieve pertinent data.
 We delve deeper into the influence of primary elements within the context memory.
@@
 Conduct experiments on Mistral-7B-inst-v0.2 using the default parameters with a context window length of 1024.
-Number of Representative Tokens.
-Q-LLM divides key-value vectors into memory blocks and picks a few representative tokens from each block to act as the segment’s representation.
+B.1 Number of Representative Tokens
+QLLM divides key-value vectors into memory blocks and picks a few representative tokens from each block to act as the segment’s representation.
 The capacity of these tokens to symbolize the entire segment semantically directly impacts the model’s efficacy.
@@
 Future work could enhance model performance by developing more efficient and potent segment representations.
-Memory Block Size.
+B.2 Memory Block Size
 Each memory block should ideally represent a consistent semantic block.
 Oversized blocks can obstruct precise lookup, while undersized ones can escalate the computational cost of memory lookup.
-We test Q-LLM with different block sizes while maintaining a total context length of $1024$.
+We test QLLM with different block sizes while maintaining a total context length of $1024$.
 The outcomes are displayed in  Fig.
@@
 Thus, investigating how to dynamically segment context is a vital future research direction.
-Number of Memory blocks.
+B.3 Number of Memory blocks
 The chosen blocks are employed to supply relevant context to LLMs.
@@
 Appendix C Configuration Details
+Figure 9: Examples of our QuickLLaMA-8B summarizing multiple papers Li et al. (2023c, 2024a, a, 2024b).
 Context Window Local Tokens Block Size Block Num
@@
 Table 3: The parameters for different context window length, including number of local tokens, memory block size and number of memory blocks.
-Datasets. We utilize representative tasks from following widely-recognized long document benchmarks:
+C.1 Datasets
+We utilize representative tasks from following widely-recognized long document benchmarks.
 (1) $\infty$-Bench Zhang et al.
@@
 This test requires the model to accurately reproduce the details from a specific sentence (referred to as the needle) that is randomly positioned within a document that could be as long as 128K (referred to as the haystack).
-Baseline Methods.
+We adopted the following setting: the needle is The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day., and the haystack is PaulGrahamEssays.
+The retrieval question is What is the best thing to do in San Francisco?.
+(3) Scaling up.
+To evaluate QLLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K.
+We use the Retrieve.KV task from the $\infty$-Bench to test its ability to discern context in extensive sequences.
+This task requires LLMs to identify a value from a key and a dictionary, essentially locating pertinent information within long sequences.
+Inputs with $\{32,64,128,256,768,1024\}$ thousand tokens are automatically generated.
+For each length, 50 instances are created for assessment.
+(4) BABILong is a challenging benchmark for evaluating the performance of models in processing arbitrarily long documents with distributed facts.
+BABI tasks are generated by simulating a set of characters and objects engaged in various movements and interactions with each other in multiple locations.
+Each interaction is represented by a fact, and the task is to answer a question using the facts from the current simulation.
+The bAbI tasks vary based on the number of facts, question complexity and the aspects of reasoning.
+C.2 Baseline Methods
 Our goal is to enable LLMs trained with limited sequence lengths to comprehend extremely long sequences without additional training.
@@
 (2024a).
-The same prompts are utilized for all methods and detailed in Appendix F.
-Implementation Details.
+The same prompts are utilized for all methods and detailed in Appendix E.
+C.3 Implementation Details
 Following the settings of previous works Xiao et al.
 (2024a), the number of representative tokens, $r_{k}$, is set to 4.
-The number of initial tokens is set to 128 for LM-Infinite, StreamingLLM, InfLLM, and Q-LLM, to cover system prompts and task descriptions.
+The number of initial tokens is set to 128 for LM-Infinite, StreamingLLM, InfLLM, and QLLM, to cover system prompts and task descriptions.
 We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K).
-The local tokens, block size, and number of blocks for InfLLM and Q-LLM are detailed in Tab.
+The local tokens, block size, and number of blocks for InfLLM and QLLM are detailed in Tab.
 3.
-Appendix D Detailed Related Works
-The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al.
-(2023); Tay et al.
-(2022); Huang et al.
-(2023); Chen et al.
-(2023b); Yu et al.
-(2023).
-The approaches can generally be categorized into two main methods, our primary focus is on the efficient context computation, which seeks to improve the computational efficiency of attention layers.
-This allows LLMs to make better use of limited resources when processing longer sequences.
-To ensure the completeness of the survey, in this section, we provide an introduction to the other method, context length extrapolation, which aims to enable LLMs trained on short sequences to handle much longer sequences, as well as the relevant work on memory-based models.
-D.1 Efficient Context Computation
-The computational complexity of attention layers, which grows quadratically, is a significant bottleneck restricting LLMs’ capability to handle lengthy sequences.
-Consequently, numerous researchers have devised efficient attention mechanisms, including sparse attention Zaheer et al.
-(2021); Beltagy et al.
-(2020); Child et al.
-(2019); Ainslie et al.
-(2020); Zhao et al.
-(2019), approximate attention computations using kernel functions Kitaev et al.
-(2020); Wang et al.
-(2020); Katharopoulos et al.
-(2020), and replacing attention layers with state-space models of linear complexity Gu et al.
-(2022); Gu and Dao (2023).
-These approaches necessitate architectural modifications, requiring retraining of the models.
-Concurrently, many scholars have tackled this challenge from an infrastructural angle by optimizing the memory usage of attention computations to mitigate the computational resource requirements of the model Dao et al.
-(2022); Dao (2023); Hong et al.
-(2024); Shazeer (2019); Kwon et al.
-(2023).
-Given the training-free nature of our method, it can be seamlessly integrated to further expedite LLM inference.
-D.2 Memory-based Approaches
-Memory networks have been extensively researched for decades and have demonstrated effectiveness in enhancing models with additional information storage capabilities Graves et al.
-(2014); Weston et al.
-(2015); Sukhbaatar et al.
-(2015); Miller et al.
-(2016).
-With the rise of pre-trained models, memory layers have gradually found application in the training stage of recurrent transformer layers, enabling models to recursively process long sequences Dai et al.
-(2019); Rae et al.
-(2020); Khandelwal et al.
-(2020); Wu et al.
-(2022); Bertsch et al.
-(2023).
-These approaches segment sequences, encoding each segments individually, and utilize memory to retain context information from preceding segments.
-Although these methodologies share conceptual similarities with Q-LLM, they necessitate architectural modifications and are typically incorporated during the pre-training phase.
-In contrast, our objective is to explore the intrinsic properties of LLMs and introduce a training-free Query-aware Context Lookup mechanism for long-text comprehension.
-Figure 9: Examples of our QuickLLaMA-8B summarizing multiple papers Li et al. (2023c, 2024a, a, 2024b).
-Appendix E More Examples
+Appendix D More Examples
 In this section, we provide examples of our QuickLLaMA-8B.
@@
 Figure 12: Examples of our QuickLLaMA-8B reading long novel Journey to the West comparing with current SOTAs
-Appendix F Prompts
+Appendix E Prompts
 We utilize the same prompts for all experiments in the main paper.